### v1

# Improved GUI Grounding via Iterative Narrowing

Abstract
GUI grounding, the task of identifying a precise location on an interface image from a natural language query, plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework called Iterative Narrowing (IN) to further enhance the performance of both general and fine-tuned models in GUI grounding. For evaluation, we tested our method on a comprehensive benchmark comprising different UI platforms.
1 Introduction
GUI Grounding is the task of identifying the visual location on an interface image given a natural language query [4, 8]. This task is essential for improving the capabilities of Vision Language Model (VLM) agents, enabling them to better understand and interact with graphical user interfaces [9]. While general VLMs, such as GPT-4V, have demonstrated high performance across a variety of visual-linguistic tasks, their accuracy in GUI grounding tasks remains suboptimal [4, 9].
Existing solutions have attempted to bridge this performance gap by fine-tuning VLMs for GUI grounding, resulting in improvements over baseline models [5, 4, 8, 6]. However, these approaches often come with the need for extensive data and re-training.
Instead of training a model, we propose a visual prompting framework called Iterative Narrowing (IN) to improve the performance of current ones. Our approach treats the model’s initial position predictions as an approximation. The framework iteratively refines the prediction by focusing on progressively smaller, cropped regions that center around the previous prediction. This refinement process can be repeated for multiple iterations.
Figure 1: A visual demonstration of IN. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
We evaluate the performance of our method on the ScreenSpot [4] benchmark. Our findings demonstrate that this iterative approach leads to substantial performance gains, providing a straightforward yet effective solution for enhancing GUI grounding capabilities.
To encourage reproducibility and facilitate further research, we provide our code at https://github.com/ant-8/GUI-Grounding-via-Iterative-Narrowing.
2 Related Work
2.1 Grounding Models
Cheng et al. [4] trained SeeClick, a GUI-specific grounding model and introduced the ScreenSpot benchmark, showing that grounding pre-training enhances task performance. Gou et al. [5] expanded on this by training a pixel-level model on a dataset of 10 million GUI elements, achieving up to 20% accuracy improvements and demonstrating effective human-like visual grounding. Wu et al. [8] introduced a cross-platform dataset of 13 million GUI elements and trained two models: OS-Atlas-Base-4B and OS-Atlas-Base-7B, improving performance across diverse benchmarks.
2.2 Iterative Visual Reasoning Beyond Convolutions
Chen et al. [1] proposed an iterative visual reasoning framework that extends beyond traditional CNNs by incorporating a local module with spatial memory and a global graph-reasoning module for spatial and semantic interactions. The global module builds a graph connecting knowledge, image regions, and class assignments, enabling iterative updates through cross-feeding between the modules to refine predictions.
2.3 GUI Agents
Xie et al. [9] introduced the OSWorld benchmark, designed to evaluate an agent’s ability to complete tasks within GUI environments. One of the key categories tested is the screenshot-only scenario, where the model is provided solely with an image of the screen, without any auxiliary information. This scenario emphasizes the importance of strong grounding capabilities, as a common failure case observed by the authors stems from inadequate grounding performance.
Wu et al. [8] expanded on this challenge by integrating their pre-trained grounding model, OS-Atlas-Base-7B, with GPT-4o, demonstrating significant improvements over GPT-4o alone on the OSWorld benchmark. Their results further validate the crucial role of effective grounding in improving task completion in GUI environments.
Hong et al. [6] also addressed GUI grounding by training and evaluating CogAgent on the screenshot-only category of OSWorld. CogAgent was developed to function both as a grounding model as well as a standalone visual agent.
3 Methodology
Given an input image and a corresponding text query, our method begins by preprocessing the image to standardize its size to $9999\times 9$ 999 pixels. The model is then prompted to predict normalized coordinates $(x,y)$ that identify a relevant location on the image. Using these predicted coordinates, we create an image crop centered around the prediction for the next iteration (see Figure 1). If the original image has a landscape orientation, the new crop has its width halved and its height halved upon each iteration. For portrait-oriented images, the width is reduced to 80% of its original size and the height is reduced to one-third of the original size.
This iterative process of generating new coordinates and refining the image crop continues for $n$ iterations. On the final iteration, we take the coordinates generated as the model’s prediction of the target location after converting them to be relative to the original image.
4 Experiments
4.1 Evaluation Results
We evaluated our method on the ScreenSpot [4] benchmark, a comprehensive assessment tool for single-step GUI grounding across multiple platforms. The ScreenSpot benchmark is structured into three primary categories: mobile, web, and desktop, each designed to reflect common user interface environments. Furthermore, each category is divided into two subcategories based on the type of target element: text or icon.
For the entire evaluation, we used $n=3$ iterations. While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken. Future work may explore the impact of varying $n$ and the extent of each cropping step on performance and inference time.
Models Baseline IN
InternVL-2-4B 4.32 6.53
Qwen2-VL-7B 42.89 69.1
OS-Atlas-Base-7B 82.47 83.33
Table 1: Overall average accuracy comparing baseline against our method (IN) on the ScreenSpot benchmark.
Models Text Icon/Widget
Baseline IN Baseline IN
InternVL-2-4B 9.16 14.65 1.64 2.18
Qwen2-VL-7B 61.34 83.52 39.29 57.21
OS-Atlas-Base-7B 93.04 93.04 72.93 74.24
Table 2: Accuracy scores in the mobile category.
Models Text Icon/Widget
Baseline IN Baseline IN
InternVL-2-4B 4.64 8.76 4.29 2.14
Qwen2-VL-7B 52.01 84.54 44.98 60.71
OS-Atlas-Base-7B 91.75 92.27 62.86 77.14
Table 3: Accuracy scores in the desktop category.
Models Text Icon/Widget
Baseline IN Baseline IN
InternVL-2-4B 0.87 5.22 0.10 2.91
Qwen2-VL-7B 33.04 73.04 21.84 50.0
OS-Atlas-Base-7B 90.87 86.96 74.27 72.33
Table 4: Accuracy scores in the web category.
4.2 Weaknesses & Observations
A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues. GUI-specific models like OS-Atlas-Base-7B [8] can accurately identify such targets in the baseline, while our method often fails due to limited context.
For example, in a table screenshot with text labels on the left and action buttons on the right, our method may focus on a cropped area containing only buttons ($\texttt{A}_{\texttt{button}},\texttt{B}_{\texttt{button}},\texttt{C}_{\texttt{button}}$), losing sight of the labels needed to correctly identify the target.
Figure 2: An example of a context-dependent failure case. At iteration $k<n$ (top image), the cropping leads to the loss of crucial contextual information—specifically, the association between each star button and its corresponding project. Thus, resulting in the wrong UI element being selected (bottom image).
Our method improves precision but at the cost of underperforming in context-dependent scenarios. This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [7]) versus OS-Atlas-Base-7B [8], which already has strong precision abilities from training. Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.
4.3 Preliminary Solutions and Future Work
To address the loss of contextual information, we conducted preliminary experiments where the model was provided with both the entire screenshot and the current crop during each iteration. This approach was inspired by the work of Chen et al. [1], which introduced a mechanism to leverage both local and global modules that iteratively exchange information for enhanced reasoning. Our goal was to enable the model to retain a global understanding while refining its focus through local crops.
However, during these initial experiments, we observed that the Visual Language Model (VLM) frequently confused the local crop with the global context image. This confusion led to incorrect coordinate predictions, as the model sometimes generated coordinates referencing the entire global image rather than accurately focusing on the local context intended for refinement.
We believe that further fine-tuning could help the model more effectively distinguish between global and local contexts, ultimately improving its performance in context-dependent scenarios.
5 Conclusion
In this work, we introduced Iterative Narrowing (IN), a visual prompting framework designed to enhance the GUI grounding capabilities of Vision-Language Models (VLMs). Through iterative refinement of predictions, IN improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface. Evaluation on the ScreenSpot [4] benchmark demonstrated that IN improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [7], where substantial improvements were observed compared to baselines. However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.
Future work may focus on addressing these contextual limitations by incorporating both global and local context information more effectively. Promising directions may include refining the model’s ability to differentiate between local crops and the entire image. There is potential to further push the boundaries of VLM precision in GUI grounding tasks, contributing to the development of more effective visual agents.

### v2

@@
 Abstract
-GUI grounding, the task of identifying a precise location on an interface image from a natural language query, plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents.
+Graphical User Interface (GUI) grounding plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents.
 While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal.
 Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance.
-We introduce a visual prompting framework called Iterative Narrowing (IN) to further enhance the performance of both general and fine-tuned models in GUI grounding.
-For evaluation, we tested our method on a comprehensive benchmark comprising different UI platforms.
+We introduce a visual prompting framework that employs an iterative narrowing mechanism to improve the performance of both general and fine-tuned models in GUI grounding by up to 61%.
+For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
 1 Introduction
@@
 However, these approaches often come with the need for extensive data and re-training.
-Instead of training a model, we propose a visual prompting framework called Iterative Narrowing (IN) to improve the performance of current ones.
+Instead of training a model, we propose a visual prompting framework that employs an iterative narrowing mechanism to improve the grounding performance of current VLMs.
 Our approach treats the model’s initial position predictions as an approximation.
@@
 This refinement process can be repeated for multiple iterations.
-Figure 1: A visual demonstration of IN. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
+Figure 1: A visual demonstration of our method. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
 We evaluate the performance of our method on the ScreenSpot [4] benchmark.
@@
 [9] introduced a cross-platform dataset of 13 million GUI elements and trained two models: OS-Atlas-Base-4B and OS-Atlas-Base-7B, improving performance across diverse benchmarks.
-2.2 Iterative Visual Reasoning Beyond Convolutions
-Chen et al.
-[1] proposed an iterative visual reasoning framework that extends beyond traditional CNNs by incorporating a local module with spatial memory and a global graph-reasoning module for spatial and semantic interactions.
-The global module builds a graph connecting knowledge, image regions, and class assignments, enabling iterative updates through cross-feeding between the modules to refine predictions.
 2.2 GUI Agents
@@
 CogAgent was developed to function both as a grounding model as well as a standalone visual agent.
+2.3 Iterative Visual Reasoning Beyond Convolutions
+Chen et al.
+[1] proposed an iterative visual reasoning framework that extends beyond traditional CNNs by incorporating a local module with spatial memory and a global graph-reasoning module for spatial and semantic interactions.
+The global module builds a graph connecting knowledge, image regions, and class assignments, enabling iterative updates through cross-feeding between the modules to refine predictions.
+2.4 Iterative Localization in Proposal-Free Referring Expression Grounding
+Sun et al.
+[7] tackled proposal-free referring expression grounding introducing a similar mechanism called iterative shrinking, which is guided by a reinforcement learning agent to localize target objects based on query sentences.
+This iterative approach parallels our method, though Sun et al. uses reinforcement learning, while our method relies on VLM predictions to refine the target.
 3 Methodology
@@
 Furthermore, each category is divided into two subcategories based on the type of target element: text or icon.
-For the entire evaluation, we used $n=3$ iterations.
-While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken.
-Future work may explore the impact of varying $n$ and the extent of each cropping step on performance and inference time.
-Models Baseline IN
+Models Baseline Ours
 InternVL-2-4B 4.32 6.53
@@
 OS-Atlas-Base-7B 82.47 83.33
-Table 1: Overall average accuracy comparing baseline against our method (IN) on the ScreenSpot benchmark.
+Table 1: Overall average accuracy comparing baseline against our method on the ScreenSpot benchmark.
 Models Text Icon/Widget
-Baseline IN Baseline IN
+Baseline Ours Baseline Ours
 InternVL-2-4B 9.16 14.65 1.64 2.18
@@
 Models Text Icon/Widget
-Baseline IN Baseline IN
+Baseline Ours Baseline Ours
 InternVL-2-4B 4.64 8.76 4.29 2.14
@@
 Models Text Icon/Widget
-Baseline IN Baseline IN
+Baseline Ours Baseline Ours
 InternVL-2-4B 0.87 5.22 0.10 2.91
@@
 Table 4: Accuracy scores in the web category.
+For the entire evaluation, we used $n=3$ iterations.
+While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken.
+Future work may explore the impact of varying $n$ and the extent of each cropping step on performance and inference time.
 4.2 Weaknesses & Observations
@@
 5 Conclusion
-In this work, we introduced Iterative Narrowing (IN), a visual prompting framework designed to enhance the GUI grounding capabilities of Vision-Language Models (VLMs).
-Through iterative refinement of predictions, IN improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface.
-Evaluation on the ScreenSpot [4] benchmark demonstrated that IN improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [7], where substantial improvements were observed compared to baselines.
+In this work, we introduced a visual prompting framework designed to perform an iterative narrowing mechanism.
+As a result, this enhances the GUI grounding capabilities of Vision-Language Models (VLMs).
+Through iterative refinement of predictions, our method improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface.
+Evaluation on the ScreenSpot [4] benchmark demonstrated that our method improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [8], where substantial improvements were observed compared to baselines.
 However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.

### v4

@@
 Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance.
-We introduce a visual prompting framework that employs an iterative narrowing mechanism to improve the performance of both general and fine-tuned models in GUI grounding by up to 61%.
+We introduce a visual prompting framework that employs an iterative narrowing mechanism to further improve the performance of both general and fine-tuned models in GUI grounding.
 For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@
 3 Methodology
-Given an input image and a corresponding text query, our method begins by preprocessing the image to standardize its size to $9999\times 9$ 999 pixels.
-The model is then prompted to predict normalized coordinates $(x,y)$ that identify a relevant location on the image.
-Using these predicted coordinates, we create an image crop centered around the prediction for the next iteration (see Figure 1).
-If the original image has a landscape orientation, the new crop has its width halved and its height halved upon each iteration.
-For portrait-oriented images, the width is reduced to 80% of its original size and the height is reduced to one-third of the original size.
-This iterative process of generating new coordinates and refining the image crop continues for $n$ iterations.
-On the final iteration, we take the coordinates generated as the model’s prediction of the target location after converting them to be relative to the original image.
+Our method refines the prediction of target coordinates within an input image $\mathcal{I}$, guided by a text query $\mathcal{Q}$, through an iterative process of cropping and coordinate prediction.
+Initially, the entire image $\mathcal{I}$ is treated as the cropping window, with dimensions equal to the original image size.
+At each iteration, the model predicts normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$, representing a relevant location within the current cropping window.
+These coordinates are used to update the cropping window for the next iteration, progressively narrowing the region of focus.
+The cropping window dimensions are adjusted based on the image orientation.
+For landscape-oriented images, the width and height are halved at each iteration.
+In the case of portrait-oriented images, the width is scaled less aggressively by a factor of $1/1.2$, while the height is halved.
+After $n$ iterations, the final normalized coordinates are transformed into absolute coordinates relative to the original image dimensions.
+This transformation ensures that the output $(X,Y)$ corresponds directly to a location within the original image.
+3.1 Pseudocode
+The following pseudocode outlines the iterative process of predicting the target location in the image:
+Algorithm 1 Iterative Cropping and Prediction
+Input: Image $\mathcal{I}$, Query $\mathcal{Q}$, Number of iterations $n$
+Output: Final coordinates $(X,Y)$ relative to the original image
+  1. 1.
+    Initialize cropping window $\mathcal{C}_{0}$ dimensions $w_{0}=W_{\text{orig}},h_{0}=H_{\text{orig}}$
+  2. 2.
+    For $k=1$ to $n$:
+    1. (a)
+      Predict normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$ relative to $\mathcal{C}_{k-1}$, from $\mathcal{Q}$
+    2. (b)
+      If $\mathcal{I}$ is landscape-oriented:
+      • $w_{k}\leftarrow\frac{w_{k-1}}{2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
+    3. (c)
+      Else (portrait-oriented):
+      • $w_{k}\leftarrow\frac{w_{k-1}}{1.2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
+    4. (d)
+      Define cropping window $\mathcal{C}_{k}$:
+      • Center: $(x_{k},y_{k})$
+      • Dimensions: $(w_{k},h_{k})$
+  3. 3.
+    Convert normalized $(x_{n},y_{n})$ to coordinates relative to the original image space $(X,Y)$
+  4. 4.
+    Return Final coordinates $(X,Y)$
 4 Experiments
@@
 A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues.
-GUI-specific models like OS-Atlas-Base-7B [9] can accurately identify such targets in the baseline, while our method often fails due to limited context.
-For example, in a table screenshot with text labels on the left and action buttons on the right, our method may focus on a cropped area containing only buttons ($\texttt{A}_{\texttt{button}},\texttt{B}_{\texttt{button}},\texttt{C}_{\texttt{button}}$), losing sight of the labels needed to correctly identify the target.
+GUI-specific models like OS-Atlas-Base-7B [9] can accurately identify such targets in the baseline, while our method often fails due to limited context (see Figure 2).
 Figure 2: An example of a context-dependent failure case.

### v5

@@
 While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal.
-Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance.
-We introduce a visual prompting framework that employs an iterative narrowing mechanism to further improve the performance of both general and fine-tuned models in GUI grounding.
+Recent studies have focused on fine-tuning these models specifically for zero-shot GUI grounding, yielding significant improvements over baseline performance.
+We introduce a visual prompting framework that employs an iterative narrowing mechanism to further improve the performance of both base and finetuned models in GUI grounding.
 For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@
 This refinement process can be repeated for multiple iterations.
-Figure 1: A visual demonstration of our method. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
+Figure 1: A visual demonstration of our method. On each iteration, the model prediction (red cross) determines the next region (red box) to iterate.
 We evaluate the performance of our method on the ScreenSpot [4] benchmark.
@@
 Cheng et al.
-[4] trained SeeClick, a GUI-specific grounding model and introduced the ScreenSpot benchmark, showing that grounding pre-training enhances task performance.
+[4] trained SeeClick, a finetuned grounding model, and introduced the ScreenSpot benchmark, demonstrating the benefits of grounding pre-training.
 Gou et al.
-[5] expanded on this by training a pixel-level model on a dataset of 10 million GUI elements, achieving up to 20% accuracy improvements and demonstrating effective human-like visual grounding.
+[5] achieved up to 20% accuracy gains with a pixel-level model trained on 10 million GUI elements.
 Wu et al.
-[9] introduced a cross-platform dataset of 13 million GUI elements and trained two models: OS-Atlas-Base-4B and OS-Atlas-Base-7B, improving performance across diverse benchmarks.
+[10] developed OS-Atlas models using a 13-million GUI element dataset, improving performance across benchmarks.
+Lin et al.
+[7] proposed ShowUI, a lightweight vision-language-action model optimized for GUI tasks using innovative token selection and streaming methods.
 2.2 GUI Agents
 Xie et al.
-[10] introduced the OSWorld benchmark, designed to evaluate an agent’s ability to complete tasks within GUI environments.
-One of the key categories tested is the screenshot-only scenario, where the model is provided solely with an image of the screen, without any auxiliary information.
-This scenario emphasizes the importance of strong grounding capabilities, as a common failure case observed by the authors stems from inadequate grounding performance.
+[11] introduced the OSWorld benchmark to evaluate GUI task agents, emphasizing grounding in screenshot-only scenarios.
 Wu et al.
-[9] expanded on this challenge by integrating their pre-trained grounding model, OS-Atlas-Base-7B, with GPT-4o, demonstrating significant improvements over GPT-4o alone on the OSWorld benchmark.
-Their results further validate the crucial role of effective grounding in improving task completion in GUI environments.
+[10] enhanced task performance by integrating OS-Atlas-Base-7B with GPT-4o.
 Hong et al.
-[6] also addressed GUI grounding by training and evaluating CogAgent on the screenshot-only category of OSWorld.
-CogAgent was developed to function both as a grounding model as well as a standalone visual agent.
-2.3 Iterative Visual Reasoning Beyond Convolutions
+[6] addressed GUI grounding with CogAgent, designed as both a grounding model and visual agent.
+2.3 Iterative Reasoning and Localization
 Chen et al.
-[1] proposed an iterative visual reasoning framework that extends beyond traditional CNNs by incorporating a local module with spatial memory and a global graph-reasoning module for spatial and semantic interactions.
-The global module builds a graph connecting knowledge, image regions, and class assignments, enabling iterative updates through cross-feeding between the modules to refine predictions.
-2.4 Iterative Localization in Proposal-Free Referring Expression Grounding
-Sun et al.
-[7] tackled proposal-free referring expression grounding introducing a similar mechanism called iterative shrinking, which is guided by a reinforcement learning agent to localize target objects based on query sentences.
-This iterative approach parallels our method, though Sun et al. uses reinforcement learning, while our method relies on VLM predictions to refine the target.
+[1] introduced a visual reasoning framework combining spatial memory and graph reasoning modules to iteratively refine predictions through spatial and semantic interactions.
+Similarly, Sun et al.
+[8] proposed an iterative shrinking approach for referring expression grounding using reinforcement learning, which parallels our method’s reliance on VLM predictions for target refinement.
 3 Methodology
@@
 These coordinates are used to update the cropping window for the next iteration, progressively narrowing the region of focus.
-The cropping window dimensions are adjusted based on the image orientation.
-For landscape-oriented images, the width and height are halved at each iteration.
-In the case of portrait-oriented images, the width is scaled less aggressively by a factor of $1/1.2$, while the height is halved.
+The cropping window dimensions are adjusted dynamically based on the image orientation.
+For landscape-oriented images, the width and height are scaled by factors $\alpha_{\text{landscape}}$ and $\beta_{\text{landscape}}$, respectively, where both scale factors are less than one.
+Similarly, for portrait-oriented images, the width is scaled by $\alpha_{\text{portrait}}$, which is less aggressive than the landscape scaling, while the height is scaled by $\beta_{\text{portrait}}$.
 After $n$ iterations, the final normalized coordinates are transformed into absolute coordinates relative to the original image dimensions.
@@
     1. (a)
-      Predict normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$ relative to $\mathcal{C}_{k-1}$, from $\mathcal{Q}$
+      Using VLM: Predict normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$ relative to $\mathcal{C}_{k-1}$, from $\mathcal{Q}$
     2. (b)
       If $\mathcal{I}$ is landscape-oriented:
-      • $w_{k}\leftarrow\frac{w_{k-1}}{2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
+      • $w_{k}\leftarrow\alpha_{\text{landscape}}\cdot w_{k-1},\quad h_{k}\leftarrow\beta_{\text{landscape}}\cdot h_{k-1}$
     3. (c)
       Else (portrait-oriented):
-      • $w_{k}\leftarrow\frac{w_{k-1}}{1.2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
+      • $w_{k}\leftarrow\alpha_{\text{portrait}}\cdot w_{k-1},\quad h_{k}\leftarrow\beta_{\text{portrait}}\cdot h_{k-1}$
     4. (d)
@@
 4 Experiments
-4.1 Evaluation Results
+4.1 Setup
 We evaluated our method on the ScreenSpot [4] benchmark, a comprehensive assessment tool for single-step GUI grounding across multiple platforms.
@@
 Furthermore, each category is divided into two subcategories based on the type of target element: text or icon.
-Models Baseline Ours
-InternVL-2-4B 4.32 6.53
+For the entire evaluation, we used $n=3$ iterations.
+While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken.
+The cropping aggressiveness was controlled using scaling factors:
+  • For landscape-oriented images:
+    • Width scaling factor: $\alpha_{\text{landscape}}=0.5$
+    • Height scaling factor: $\beta_{\text{landscape}}=0.5$
+  • For portrait-oriented images:
+    • Width scaling factor: $\alpha_{\text{portrait}}=\frac{1}{1.2}\approx 0.833$
+    • Height scaling factor: $\beta_{\text{portrait}}=0.5$
+Future work may explore the impact of varying $n$ and the values of $\alpha_{\text{landscape}}$, $\beta_{\text{landscape}}$, $\alpha_{\text{portrait}}$, and $\beta_{\text{portrait}}$ on performance and inference time.
+4.2 Evaluation Results
+Model Type Models Baseline IN
+Base InternVL-2-4B 4.32 6.53
 Qwen2-VL-7B 42.89 69.1
+Finetuned ShowUI-2B 75.1 79.56
 OS-Atlas-Base-7B 82.47 83.33
-Table 1: Overall average accuracy comparing baseline against our method on the ScreenSpot benchmark.
-Models Text Icon/Widget
-Baseline Ours Baseline Ours
-InternVL-2-4B 9.16 14.65 1.64 2.18
+Table 1: Overall average accuracy (%) comparing baseline against our method (IN) on the ScreenSpot benchmark.
+Model Type Models Text Icon/Widget
+Baseline IN Baseline IN
+Base InternVL-2-4B 9.16 14.65 1.64 2.18
 Qwen2-VL-7B 61.34 83.52 39.29 57.21
+Finetuned ShowUI-2B 92.3 89.01 75.5 74.24
 OS-Atlas-Base-7B 93.04 93.04 72.93 74.24
 Table 2: Accuracy scores in the mobile category.
-Models Text Icon/Widget
-Baseline Ours Baseline Ours
-InternVL-2-4B 4.64 8.76 4.29 2.14
+Model Type Models Text Icon/Widget
+Baseline IN Baseline IN
+Base InternVL-2-4B 4.64 8.76 4.29 2.14
 Qwen2-VL-7B 52.01 84.54 44.98 60.71
+Finetuned ShowUI-2B 76.3 87.6 61.1 68.57
 OS-Atlas-Base-7B 91.75 92.27 62.86 77.14
 Table 3: Accuracy scores in the desktop category.
-Models Text Icon/Widget
-Baseline Ours Baseline Ours
-InternVL-2-4B 0.87 5.22 0.10 2.91
+Model Type Models Text Icon/Widget
+Baseline IN Baseline IN
+Base InternVL-2-4B 0.87 5.22 0.10 2.91
 Qwen2-VL-7B 33.04 73.04 21.84 50.0
+Finetuned ShowUI-2B 81.7 83.04 63.6 68.93
 OS-Atlas-Base-7B 90.87 86.96 74.27 72.33
 Table 4: Accuracy scores in the web category.
-For the entire evaluation, we used $n=3$ iterations.
-While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken.
-Future work may explore the impact of varying $n$ and the extent of each cropping step on performance and inference time.
 4.3 Weaknesses & Observations
-A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues.
-GUI-specific models like OS-Atlas-Base-7B [9] can accurately identify such targets in the baseline, while our method often fails due to limited context (see Figure 2).
+A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues (see Figure 2).
 Figure 2: An example of a context-dependent failure case.
@@
 Our method improves precision but at the cost of underperforming in context-dependent scenarios.
-This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [8]) versus OS-Atlas-Base-7B [9], which already has strong precision abilities from training.
+This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [9]) versus OS-Atlas [10] and ShowUI [7], which already has strong precision abilities from training.
 Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.

### v6

@@
 1 Introduction
-GUI Grounding is the task of identifying the visual location on an interface image given a natural language query [4, 10].
-This task is essential for improving the capabilities of Vision Language Model (VLM) agents, enabling them to better understand and interact with graphical user interfaces [11].
-While general VLMs, such as GPT-4V, have demonstrated high performance across a variety of visual-linguistic tasks, their accuracy in GUI grounding tasks remains suboptimal [4, 11].
-Existing solutions have attempted to bridge this performance gap by fine-tuning VLMs for GUI grounding, resulting in improvements over baseline models [5, 4, 10, 6].
+GUI Grounding is the task of identifying the visual location on an interface image given a natural language query Cheng et al.
+(2024); Wu et al.
+(2024).
+This task is essential for improving the capabilities of Vision Language Model (VLM) agents, enabling them to better understand and interact with graphical user interfaces Xie et al.
+(2024).
+While general VLMs, such as GPT-4V, have demonstrated high performance across a variety of visual-linguistic tasks, their accuracy in GUI grounding tasks remains suboptimal (Cheng et al., 2024; Xie et al., 2024).
+Existing solutions have attempted to bridge this performance gap by fine-tuning VLMs for GUI grounding, resulting in improvements over baseline models (Gou et al., 2024; Cheng et al., 2024; Wu et al., 2024; Hong et al., 2023).
 However, these approaches often come with the need for extensive data and re-training.
@@
 Figure 1: A visual demonstration of our method. On each iteration, the model prediction (red cross) determines the next region (red box) to iterate.
-We evaluate the performance of our method on the ScreenSpot [4] benchmark.
+We evaluate the performance of our method on the ScreenSpot (Cheng et al., 2024) benchmark.
 Our findings demonstrate that this iterative approach leads to substantial performance gains, providing a straightforward yet effective solution for enhancing GUI grounding capabilities.
@@
 2 Related Work
+2.1 Iterative Reasoning and Localization
+Iterative refinement strategies have been explored in multiple contexts, aiming to progressively narrow down the target location.
+Wu et al.
+Wu and Xie (2023) proposed SEAL, an LLM-guided visual search framework designed for multimodal large language models in high-resolution natural image understanding, particularly for visual question answering tasks.
+SEAL integrates a dedicated visual search model, specialized localization decoders, and a Visual Working Memory to actively recover missing visual details, requiring additional model inference beyond the base VLM.
+In contrast, our iterative narrowing method is developed specifically for precise GUI element grounding, is training-free, and reuses the same VLM across successive refinement steps.
+Chen et al.
+Chen et al.
+(2018) introduced a visual reasoning framework that combines spatial memory with graph reasoning to iteratively refine predictions through spatial and semantic cues.
+Similarly, Sun et al.
+Sun et al.
+(2021) proposed an iterative shrinking approach for referring expression grounding and relies on reinforcement learning to determine the shrinking strategy, whereas we perform refinement purely through repeated model predictions without additional training or policy learning.
 2.2 Grounding Models
+Recent work on GUI grounding has explored both large-scale pretraining and task-specific model design.
 Cheng et al.
-[4] trained SeeClick, a finetuned grounding model, and introduced the ScreenSpot benchmark, demonstrating the benefits of grounding pre-training.
+(Cheng et al., 2024) trained SeeClick, a finetuned grounding model, and introduced the ScreenSpot benchmark, highlighting the benefits of grounding-specific pretraining.
 Gou et al.
-[5] achieved up to 20% accuracy gains with a pixel-level model trained on 10 million GUI elements.
+(Gou et al., 2024) achieved up to 20% accuracy gains using a pixel-level model trained on 10 million GUI elements.
 Wu et al.
-[10] developed OS-Atlas models using a 13-million GUI element dataset, improving performance across benchmarks.
+(Wu et al., 2024) developed OS-Atlas models with a 13-million GUI element dataset, delivering consistent improvements across multiple benchmarks.
 Lin et al.
-[7] proposed ShowUI, a lightweight vision-language-action model optimized for GUI tasks using innovative token selection and streaming methods.
+(Lin et al., 2024) proposed ShowUI, a lightweight vision-language-action model optimized for GUI tasks through efficient token selection and streaming methods.
 2.3 GUI Agents
+Beyond grounding in static scenarios, other works have explored GUI agents that perform end-to-end interaction.
 Xie et al.
-[11] introduced the OSWorld benchmark to evaluate GUI task agents, emphasizing grounding in screenshot-only scenarios.
+Xie et al.
+(2024) introduced the OSWorld benchmark to evaluate such agents, emphasizing grounding in screenshot-only settings.
 Wu et al.
-[10] enhanced task performance by integrating OS-Atlas-Base-7B with GPT-4o.
+Wu et al.
+(2024) improved task performance by combining OS-Atlas-Base-7B with GPT-4o  OpenAI (2024) for enhanced reasoning and perception.
 Hong et al.
-[6] addressed GUI grounding with CogAgent, designed as both a grounding model and visual agent.
-2.3 Iterative Reasoning and Localization
-Chen et al.
-[1] introduced a visual reasoning framework combining spatial memory and graph reasoning modules to iteratively refine predictions through spatial and semantic interactions.
-Similarly, Sun et al.
-[8] proposed an iterative shrinking approach for referring expression grounding using reinforcement learning, which parallels our method’s reliance on VLM predictions for target refinement.
+Hong et al.
+(2023) presented CogAgent, which serves as both a grounding model and an interactive visual agent.
 3 Methodology
@@
 4.1 Setup
-We evaluated our method on the ScreenSpot [4] benchmark, a comprehensive assessment tool for single-step GUI grounding across multiple platforms.
+We evaluated our method on the ScreenSpot (Cheng et al., 2024) benchmark, a comprehensive assessment tool for single-step GUI grounding across multiple platforms.
 The ScreenSpot benchmark is structured into three primary categories: mobile, web, and desktop, each designed to reflect common user interface environments.
@@
 Figure 2: An example of a context-dependent failure case.
-At iteration $k<n$ (top image), the cropping leads to the loss of crucial contextual information—specifically, the association between each star button and its corresponding project.
+At iteration $k<n$ (top image), the cropping leads to the loss of crucial contextual information.
+Specifically, the association between each star button and its corresponding project.
 Thus, resulting in the wrong UI element being selected (bottom image).
 Our method improves precision but at the cost of underperforming in context-dependent scenarios.
-This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [9]) versus OS-Atlas [10] and ShowUI [7], which already has strong precision abilities from training.
+This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B Chen et al.
+(2024, 2023), Qwen2-VL-7B Wang et al.
+(2024)) versus OS-Atlas Wu et al.
+(2024) and ShowUI Lin et al.
+(2024), which already has strong precision abilities from training.
 Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.
-4.4 Preliminary Solutions and Future Work
+4.4 Potential Solutions and Future Work
 To address the loss of contextual information, we conducted preliminary experiments where the model was provided with both the entire screenshot and the current crop during each iteration.
-This approach was inspired by the work of Chen et al.
-[1], which introduced a mechanism to leverage both local and global modules that iteratively exchange information for enhanced reasoning.
-Our goal was to enable the model to retain a global understanding while refining its focus through local crops.
-However, during these initial experiments, we observed that the Visual Language Model (VLM) frequently confused the local crop with the global context image.
+The goal was to enable the model to retain a global understanding while refining its focus through local crops.
+However, during these initial experiments, we observed that the VLM frequently confused the local crop with the global context image.
 This confusion led to incorrect coordinate predictions, as the model sometimes generated coordinates referencing the entire global image rather than accurately focusing on the local context intended for refinement.
@@
 Through iterative refinement of predictions, our method improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface.
-Evaluation on the ScreenSpot [4] benchmark demonstrated that our method improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [9], where substantial improvements were observed compared to baselines.
+Evaluation on the ScreenSpot Cheng et al.
+(2024) benchmark demonstrated that our method improves performance, especially for generalist VLMs like InternVL-2-4B Chen et al.
+(2023, 2024) and Qwen2-VL-7B Wang et al.
+(2024), where substantial improvements were observed compared to baselines.
 However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.